Small pieces were three dollars and large pizzas for five dollars. With the throne, it was necessary to spend $475 for 125 pizzas. How many pieces of each type of purchased?
1
Expert's answer
2019-05-17T11:47:41-0400

We can solve this, using System of Equations


Let`s say that we should buy x small pizzas and y large pizzas.

Now we can create system of equations.

First equation will be


"3x+5y=475"


and the second one


"x+y=125"

We can solve this one, by substitution :


Step 1:

Solve x+y=125 for x:


"x+y=125\\
x+y+(−y)=125+(−y)\\
x=−y+125"


Step 2:

Substitute −y+125 for x in 3x+5y=475:


"3x+5y=475\\
3(−y+125)+5y=475\\
2y+375=475\\
2y+375+(−375)=475+(−375)\\
2y=100\\
2y/2=100/2\\
y=50"


Step 3:

Substitute 50 for y in x=−y+125:


"x=−y+125\\
x=−50+125\\
x=75"


Answer:

We should buy 75 small pizzas and 50 large pizzas.